OpenStudy (anonymous):

Find the position vector of a particle which satisfies a(t) = with v(0) = <1, 1> and r(0) = <2, -1>. ???

OpenStudy (anonymous):

a(t) = what?

OpenStudy (anonymous):

sorry , a(t) <sin t, e^t>

OpenStudy (anonymous):

I integrated twice a(t) to get position but when I set t = 0 it does not give me what it says there

OpenStudy (anonymous):

since a(t) = v'(t), \[v'(t) = <\sin(t),e^t>\] \[v(t) = <-\cos(t) + c_1, e^t + c_2>\] using the initial conditions, \[v(0) = <-1+c_1,1+c_2> = <1,1>\] so \[c_1=2,c_2=0\] and \[v(t) = <2-\cos(t),e^t>\] integrating again, \[r(t) = <2t - \sin(t) + c_3,e^t+c_4>\] \[r(0)=<c_3,1+c_4> = <2,-1>\] so \[c_3=2,c_4=-2\] and finally \[r(t)=<2t-\sin(t)+2,e^t-2>\]

OpenStudy (anonymous):

perfect! Thank you so much!